OpenStudy (anonymous):

multiply (x^2+y^4)(x^2-y^4)

jimthompson5910 (jim_thompson5910):

use the difference of squares rule (x^2+y^4)(x^2-y^4) = (x^2)^2 - (y^4)^2 (x^2+y^4)(x^2-y^4) = x^(2*2) - y^(4*2) (x^2+y^4)(x^2-y^4) = x^4 - y^8

OpenStudy (anonymous):

thank you